xx

Probability

February 4, 2022

Contents Start of lecture 1

Example 0. Dice: outcomes 1, 2, x, 6.

ℙ(2) = \frac{1}{6}

ℙ(multiple of 3) = \frac{2}{6} = \frac{1}{3}.

ℙ(prime or a multiple of 3)  [figure: = \frac{1}{3} + \frac{1}{2} = \frac{5}{6}]

= \frac{4}{6} = \frac{2}{3}

= \frac{1}{3} + \frac{1}{2} − ℙ(prime and a multiple of 3)

= \frac{1}{3} + \frac{1}{2} − \frac{1}{6} = \frac{2}{3}

ℙ(not a multiple of 3) = \frac{2}{3}.